[Hallc_running] [New Logentry] RC Update

ddutta at jlab.org ddutta at jlab.org
Fri Jan 12 08:50:02 EST 2018

Logentry Text:
Yesterday the Hall was locked up soon after the DOE Under-secretary's tour of the Hall. Last night over the swing and owl shifts MCC was able to configure Halls A and B and send beam to all 3 Halls (D,A, B) at 5 pass. 
This means that today during the day shift they will be setting up Hall C for  3 pass beam. If all goes well we expect to see beam in the Hall during the day shift. 
Dave Meekins is already working on condensing hydrogen in the target. 
SHMS is at 8 deg, ready to start the beam steering test as soon as beam is restored.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3511339
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180112/25d254bf/attachment.html>

More information about the Hallc_running mailing list